OpenStudy (dj3strella):

can someone help me i know my answer i just dont understand the last step to getting my answer . Use the quadratic formula to solve 2y2 – 6y – 8 = 0.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2y2 – 6y – 8 = 0 is a formula in the form of ax^2 + bx + c = 0

OpenStudy (anonymous):

\[\large x_{1,2}=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

that's the quadratic formula, you just need to fill in the a, b and c and you'll get the two answers

OpenStudy (dj3strella):